OpenStudy (anonymous):

Evaluate 8 E - n/4 n=2

OpenStudy (anonymous):

@ganeshie8 . do you know what this is ?

ganeshie8 (ganeshie8):

?

ganeshie8 (ganeshie8):

You wanto evaluate \(8E(-n/4)\) when n=2 ha ?

OpenStudy (anonymous):

sorry n/5

ganeshie8 (ganeshie8):

easy, is that -n/4 or -n/5 ?

OpenStudy (anonymous):

-n/5

OpenStudy (anonymous):

\[\sum_{n=2}^{8} -\frac{ n }{ 5 }\]

ganeshie8 (ganeshie8):

\(\sum_{n=2}^{8} - \frac{ n }{ 5 } \)

ganeshie8 (ganeshie8):

uhh... why i am not getting n=2 in bottom os SUM notation

OpenStudy (anonymous):

im not sure. :(

ganeshie8 (ganeshie8):

ok, its called SUMMATION

ganeshie8 (ganeshie8):

∑ this symbol is called summation symbol

OpenStudy (anonymous):

yes. i know .

ganeshie8 (ganeshie8):

what it says is, you need to take sum. il show a simple example

OpenStudy (anonymous):

okk

ganeshie8 (ganeshie8):

\(\large \sum_{n=1}^{3} n \)

ganeshie8 (ganeshie8):

\(\large \sum_{n=1}^{3} n = 1 + 2 + 3\)

ganeshie8 (ganeshie8):

thats one example. you just need to take sum, replacing n wid 1, 2, 3 each time

OpenStudy (anonymous):

What a sigma notation tells is that we add a series of number which add up to something. So in your case it is asking what does -n/5 equal to when n starts at 2 and ends at 8. So you would add -(2/5)-(3/5)-(4/5)-(5/5)-(6/5)-(7/5)-(8/5)= ? Add those up and the sum is your answer.

ganeshie8 (ganeshie8):

use that example, and see if u can solv ethe present problem

OpenStudy (anonymous):

wait so i add 2 , 3, 4, 5, 6,7,8 ?

ganeshie8 (ganeshie8):

you're not given n, you're given -n/5

ganeshie8 (ganeshie8):

\(\sum_{n=2}^{8} -\frac{ n }{ 5 } = \frac{-2}{5} + \frac{-3}{5} +\frac{-4}{5} +\frac{-5}{5} +\frac{-6}{5} +\frac{-7}{5} +\frac{-8}{5} \) = ?

OpenStudy (anonymous):

ohhh okk . sooo -25/n ?

ganeshie8 (ganeshie8):

im getting -35/5

ganeshie8 (ganeshie8):

which equals -7

OpenStudy (anonymous):

oopss i counted wrong. okk yess i got -35/5=-7

ganeshie8 (ganeshie8):

ha

OpenStudy (anonymous):

is that all ?

ganeshie8 (ganeshie8):

thats all to SUMMATION or SIGMA notation.

ganeshie8 (ganeshie8):

when u see that symbol, just add repeatedly... like above

OpenStudy (anonymous):

ohh okk (: . so my answer is -7 ?

ganeshie8 (ganeshie8):

correct !

OpenStudy (anonymous):

oh that was pretty simple. i going to try the next few on my own. then ill shhow you (;

ganeshie8 (ganeshie8):

sounds very good :)

OpenStudy (anonymous):

\[\sum_{n=2}^{8}-(n-5)\] Add -2/5 , 3/5, 4/5, 5/5, 6/5,7/5,8/5 = -35/5 =-7 \[\sum_{n=3}^{7}\frac{ n }{ 4 }\] Add 3/4 , 4/4, 5/4, 6/4,7/4 =25/4

ganeshie8 (ganeshie8):

in the first problem, you mean -n/5 ?

OpenStudy (anonymous):

noo i realized i wrote it wrong.

ganeshie8 (ganeshie8):

which one is correct, -(n-5) or -n/5 ?

OpenStudy (anonymous):

-(n-5)

ganeshie8 (ganeshie8):

then wat u did above is not correct. do it again from scratch

OpenStudy (anonymous):

am i supposed to use the same concept ?

ganeshie8 (ganeshie8):

yes same concept, just keep adding, replacing n each time

OpenStudy (anonymous):

ok , can you tell me what i did wrong ?

ganeshie8 (ganeshie8):

whole thing is wrong for that q

ganeshie8 (ganeshie8):

however, u did the next q perfect ! :)

ganeshie8 (ganeshie8):

its like this :-

ganeshie8 (ganeshie8):

\(\sum_{n=2}^{8}-(n-5) \\ \\ = -(2-5) - (3-5) - (4-5) - (5-5) - (6-6) - (7-7) - (8-8)\)

OpenStudy (anonymous):

ok yay for # 2 (: . now number 1. do i add now ?

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

wait would i simplify whats in the parenthesis first ? like (2-5)=-3 ?

ganeshie8 (ganeshie8):

correct ! first simplify things inside parenthesis

OpenStudy (anonymous):

ok i got -(-3)-(-2)-(-1)-0-1-2-3

ganeshie8 (ganeshie8):

it wud become :- \(-(-3) - (-2) - (-1) - (0) - (1) - (2) - (3)\)

ganeshie8 (ganeshie8):

i got the same :)

ganeshie8 (ganeshie8):

simplify completely

OpenStudy (anonymous):

0 ?

ganeshie8 (ganeshie8):

Yes !

OpenStudy (anonymous):

(:. your so smart !

ganeshie8 (ganeshie8):

noo you're smart to learn these in few minutes oly :D

OpenStudy (anonymous):

Thankyouuu (::. okk a few more before my test. Given the arithmetic sequence –6, 1, 8…, evaluate\[\sum_{n=9}^{12_{}} a _{n}\]

OpenStudy (anonymous):

how would i plug in -6,1,and 8

ganeshie8 (ganeshie8):

It is asking you to find the sum of terms from 9 to 12

ganeshie8 (ganeshie8):

this is the sequence : –6, 1, 8....

ganeshie8 (ganeshie8):

first term, a = -6 common difference, d = ?

OpenStudy (anonymous):

+7 ? that how it goes from -6 to 1 to 8 .

ganeshie8 (ganeshie8):

thats right !

OpenStudy (anonymous):

ok (:. so whats the next step ?

ganeshie8 (ganeshie8):

\(\sum_{n=9}^{12_{}} a _{n} = a_9 + a_{10} + a_{11} + a_{12}\)

ganeshie8 (ganeshie8):

find \(a_9, a_{10}, a_{11}, \ \text{and} \ a_{12}\) of the sequence

ganeshie8 (ganeshie8):

\(a_9 = -6 + (9-1) 7 = -6 + 56 = 50\)

OpenStudy (anonymous):

ohhh ok. what formula did you use ?

ganeshie8 (ganeshie8):

since common difference is 7, \(a_{10} = 50+7 = 57\) \(a_{11} = 57+7 = 64\) \(a_{12} = 64+7 = 71\)

ganeshie8 (ganeshie8):

add all these together. (il give u the formula in the end)

OpenStudy (anonymous):

add 57+64+71 ?

ganeshie8 (ganeshie8):

dont forget \(a_9\)

OpenStudy (anonymous):

sooo 50+57+64+71 = 242

ganeshie8 (ganeshie8):

thats right !

ganeshie8 (ganeshie8):

Now getting back to the formula used when doing below :- \( a_9 = -6 + (9-1) 7 = -6 + 56 = 50 \)

ganeshie8 (ganeshie8):

Ive used the formula for "nth term in arithmetic sequence" : \(a_n = a + (n-1)d\) a = first term d = common difference

OpenStudy (anonymous):

ohh okk (:. so 242 was the answer. ? ok im going to attemt the next one on myh own. brb (:

ganeshie8 (ganeshie8):

good luck ! il be right here :)

ganeshie8 (ganeshie8):

and yes 242 is the answer

OpenStudy (anonymous):

Given the arithmetic sequence 23, 17, 11 …, evaluate \[\sum_{n=4}^{9} a _{n}\] a4+a5+a6+a7+a8+a9 a4=23+(4-1)-6=23+3(-6)=23+-18=5 a5=5+-6=-1 a6= -1+-6=-7 a7=-7+-6=-13 a8=-13+-6=-19 a9=-19+-6=-25 5+-1+-7+-13+-19+-25=-60

ganeshie8 (ganeshie8):

Looks perfect !

ganeshie8 (ganeshie8):

u knw, theres another way to do this

ganeshie8 (ganeshie8):

we will try it next time ok

OpenStudy (anonymous):

okkk (:. Using sigma notation, describe the total length of 7 pictures hung side-by-side. Then find the total length the pictures occupy along the wall if the first picture has a length of 10 inches and the length of each successive picture is 4 inches longer than the previous one.

OpenStudy (anonymous):

is sigma notation what we were doing ?

ganeshie8 (ganeshie8):

Yes. since we need to add 7 pictures, it wud be :- \(\sum_{1}^7 a_n \)

ganeshie8 (ganeshie8):

then, it tells us that, first pic has a length of 10 inches, that means, a = 10

ganeshie8 (ganeshie8):

i mean, \(a_1 = 10\)

ganeshie8 (ganeshie8):

since each successive pic is 4 inches extra, d = 4

ganeshie8 (ganeshie8):

so, the sequence of lengths wud be :- 10, 14, 18....

ganeshie8 (ganeshie8):

and there wud be 7 pics, right ?

ganeshie8 (ganeshie8):

So, n = 7

ganeshie8 (ganeshie8):

\(\sum_{1}^7 a_n = a_1 + a_2 + ...... + a_7\)

OpenStudy (anonymous):

so i would start of with a10, a14,a18 right ?